[Brahms-l] Fw: Shift report 20050124 16:00-24:00

From: flemming videbaek <videbaek@rcf.rhic.bnl.gov>
Date: Tue Jan 25 2005 - 06:44:47 EST
This ought to have worked, but apperently did not despite several succesful
tests.

----------------------------------------------------------------
Flemming Videbaek
Physics Department
Brookhaven National Laboratory

e-mail: videbaek@bnl.gov
phone: 631-344-4106
----- Original Message ----- 
From: "Apache" <apache@brahms-web.brahms.bnl.gov>
To: <brahms-l@lists.bnl.gov>; <hito@bnl.gov>; <videbaek@bnl.gov>
Sent: Monday, January 24, 2005 11:59 PM
Subject: Shift report 20050124 16:00-24:00


>
> ------------------------------------------------------------
> Date: Mon Jan 24 23:59:01 2005, Shift: 16:00-24:00
> Supervisor: CR, with: OR
> Log book#16 pages 1 - 1
>
> Shift Summary:
> ------------------------------------------------------------
> 1. Beam was stored at 20:30. MCR tried to use RF storage, but lost 1/3 of
the beam when RF stations tripped. Still, the beam was kept until the
problem was solved, at 23:00. Now, they are eager for a new physics store.
>
> 2. Because low collision rates, we took low field setting data: run
 12473 - Fs: 8@B608 (60k tr.2) , Mrs: 90@700B (71k tr.3). Took  also
pedestal runs, no. 12471 and 12472.
>
> 3. No detector problems found.
> ------------------------------------------------------------
>


_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Tue Jan 25 06:40:53 2005

This archive was generated by hypermail 2.1.8 : Tue Jan 25 2005 - 06:41:11 EST